A 1 meter tube, closed at one end, is immersed in water completely. What is the open air pressure if the water rises 9 cm in the tube? Assume temperature remains constant.
1
Expert's answer
2018-06-01T10:03:08-0400
According to the Boyle’s Law
p_1 V_1=p_2 V_2
We have
p_a V_1=pV_2 → p=(p_a V_1)/V_2 =p_a (S∙h)/(S∙(h-∆h) )=p_a h/(h-∆h)=101325∙1/(1-0.09)=111346 Pa